OpenStudy (anonymous):

Help please?? Solve 2x2 + 3x + 6 = 0. Round solutions to the nearest hundredth. (4 points) x ≈ -2.64 and x ≈ 1.14 x ≈ -1.14 and x ≈ 2.64 No real solutions x ≈ -1.11 and x ≈ -4.89

OpenStudy (sebaxtiangb):

the 2x2 is suppose to be 2x^2? or is it a multiplication?

OpenStudy (anonymous):

yes it is supposed to be 2x^2

OpenStudy (sebaxtiangb):

first solve the exponent first

OpenStudy (sebaxtiangb):

that will be 4x

OpenStudy (sebaxtiangb):

then add the 4x and the 3x

OpenStudy (anonymous):

7x

OpenStudy (sebaxtiangb):

yes

OpenStudy (lynfran):

that will not be 4x??

OpenStudy (sebaxtiangb):

then the equation is 7x + 3 = 0

OpenStudy (sebaxtiangb):

no bc u multiply \[2^2\]

OpenStudy (sebaxtiangb):

which is 4

OpenStudy (sebaxtiangb):

then add the variables which is 3x and 4x which is 7x and the equation is then 7x + 3 = 0

OpenStudy (sebaxtiangb):

then u can solve

OpenStudy (anonymous):

still confused and idk why

OpenStudy (sebaxtiangb):

what happen?

OpenStudy (anonymous):

well i added the variables and got the 7x+3=0

OpenStudy (anonymous):

but not sure what comes after

OpenStudy (sebaxtiangb):

subtract 3 from both sides

OpenStudy (anonymous):

7x=-3

OpenStudy (sebaxtiangb):